during shooting of a super hit film ,mard amitabh was waiting for his beloved amrita with
his dog.when he saw her approaching,dog was
exited and dashed to her,then back to his master
and so on,never stopping.how far would you estimate the dog ran if its speed is 30km/hr and each of them walked at 4km/hr;starting
400 m apart?
1
Expert's answer
2012-05-18T07:57:45-0400
4km/hr = 4000m/hr
They will meet in:
t = S / V, where S = 400m, V = V1 +
V2
t = 400 / (4000 + 4000) = 0.05hrs
The distance made by the dog is:
D
= t * Vd, where Vd = 30km/hr
D = 0.05 * 30 = 1.5km

Answer: 1.5km.